OpenStudy (ineedhelpnow):

how do i find the slope of the line passing through (5,-6) and (6,-5)?

OpenStudy (anonymous):

\[m = \frac{y_2 -y_1}{x_2-x_1}\]

OpenStudy (anonymous):

\[(x_1,y_1)=(5,-6)\]\[(x_2,y_2)=(6,-5)\]

OpenStudy (anonymous):

\[m=\frac{-5 +6}{6-5}\]

OpenStudy (anonymous):

m=1

OpenStudy (ineedhelpnow):

okay. i get it. i need to remember that! could you help me with writing the point-slope for and the slope intercept form of the equation of the line passing through (-1,60 with the slope of -4?